Please help and explain
Uul2t Chapter 1 Review Quiz This Quiz: 20 pt This Question: 1 pt a, what is the exact value that is 27% of 10187 The exact value is L
Lauren Maxwell Statistics Online (Spring 2019) Submit C Quiz: Chapter 1 Review Quiz his Question: 1 pt 14 of 20 (19 complete) This Quiz: 20 pts pos Question Help A polling company reported that 27% of 1018 s eyed adults said that rising gas prices are very harmful Complete parts a through d) below. b. Could the result from part (a) be the actual number of adults who said that rising gas prices are very harmful? Why or why not? O A. Yes, the resulit from part (a) could be the actual number of adults who said that rising gas prices are very harmful because the results are statistically significant O B. No, the result from part (a) could not be the actual number of adults who said that rising gas prices are very harmful because a count of people O c No the result from part (a) could not be the actual number of adults who said that rising gas prices are very harmful because that is a must result in a whole number. opinion. Click to select your answer(s). 1/2
Lauren Maxwell Statistics Online (Spring 2019) Submit C Quiz: Chapter 1 Review Quiz This Question: 1 pt 14 of 20 (19 complete) ▼ This Quiz: 20 pts poss t= Question Help polling company rep ted that 27% of 1018 surveyed adults said that nsing gas prices are very harmful Complete parts a through d) below. c. What could be the actual number of adults who said that rising gas prices are very harmful? The actual number of adults with this opinion could be(Type an integer or a decimal.) d. Among the 1018 respondents, 306 said that rising gas prices are not at all harmful. What percentage of respondents said that rising gas prices are (Round to two decimal places as needed.) 11:38 A 1/27/20
Answer #1

(a) The exact value = 0.27 * 1018 = 274.86

(b) The correct answer is option (B) as 274.86 is not a whole number.

(c) The actual number of adults with this opinion can be

275 as rounding off 274.86 to nearest whole number results in 275

(d) Percentage of respondents who said that rising gas prices are "not at all harmful" = 306/1018 * 100%

= 30.06 %